find equation of a parabola with vertex at(-2,3)and passing through (1,21)
Answer by scott8148(6628)

the general vertex equation is ___ y = a(x - h)^2 + k ; the vertex is (h,k)

21 = a(1 - -2)^2 + 3 ___ 18 = a(3)^2 ___ a = 2

y = 2(x + 2)^2 + 3